OpenStudy (anonymous):

If a sample of oxygen occupies a volume of 2.15 L at a pressure of 58.0 kPa and a temperature of 25°C, what volume would this sample occupy at 101.3 kPa and 0°C? A. 1.35 L B. 1.13 L C. 4.10 L D. 3.44 L

OpenStudy (aaronq):

use the ideal gas law: PV=nRT, when we move all the variables to one side, we get a constant \(\dfrac{PV}{nRT}=C\). when the conditions change, we can use this relation \(system_1=system_2\), so \(\dfrac{P_1V_1}{n_1R_1T_1}=\dfrac{P_2V_2}{n_2R_2T_2}\) since it's a closed system (assumed), n is constant. R is also a constant. the formula is reduced to: \(\dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}\) now just plug in your values and solve for the new volume.

OpenStudy (anonymous):

I got that, it's just that when I go to solve the equation, I end up dividing by 0. When I do that, I get 0 as my final answer.

OpenStudy (aaronq):

you need to use Kelvin